# Online Graph Learning via Time-Vertex Adaptive Filters: From Theory to Cardiac Fibrillation

### v1

@@ IV-B Updating $\mathbf{W}_{t}$
The next step of our algorithm involves recovering the GSO, $\hat{\mathbf{W}}_{t}$, from $\hat{\boldsymbol{\Psi}}_{t}$ through either of two paths. Path 2 provides a simple but biased estimate by setting $\hat{\mathbf{W}}_{t}=\hat{\boldsymbol{\Psi}}_{1,t}$. Path 1, which we detail below, optimises (9) to obtain unbiased estimates of $\hat{\mathbf{W}}_{t}$ at the expense of computational complexity.
@@
where $\beta_{t}$ is the stepsize, and $\mu_{1,t}$ is the sparsity parameter from the first sub-problem. Then, $\hat{\mathbf{W}}_{t}={\hat{\mathbf{W}}}_{+_{t}}-{\hat{\mathbf{W}}}_{-_{t}}$.
@@ IV-C Estimating $\mathbf{h}_{t}$
To fully identify the CGP model in (5), the graph filter coefficients $\hat{\mathbf{h}}_{t}$ remain to be estimated. This phase is optional since the primary objective of this work is to estimate the GSO structure. Notice that, $\hat{\mathbf{W}}_{t}$ must be debiased before estimating $\hat{\mathbf{h}}_{t}$, due to the heavy regularisation involved. Debiasing is performed by fixing and optimising only the non-zero elements of $\hat{\mathbf{W}}_{t}$ using least squares. Note that reducing the sample size in this way can risk distorting the noise distribution from normality, potentially compromising the minimum MSE criterion for noisy or small datasets [35].
After debiasing $\hat{\mathbf{W}}_{t}$, the filter coefficients $\hat{\mathbf{h}}_{t}$ are estimated. Unlike zeros in the GSO, which indicate absence of edges and thus have causal significance, sparsity in $\hat{\mathbf{h}}$ serves mainly to simplify the model structure. Given the more relaxed sparsity requirements, GAR-LMS [33] is employed to arrive at the recursive update equation of $\hat{\mathbf{h}}$, given by
@@
where $\mathbf{Y}_{t}$ contains the lagged graph shifted signals as in (11), $\mathbf{C}_{t}\in\mathbb{R}^{M\times M}$ tracks correlations between the shifted signals, $\mathbf{u}_{t}\in\mathbb{R}^{M}$ accumulates cross-correlations with the desired response, $\mathbf{b}_{t}$ is the sparsity-promoting re-weighting vector, $\rho_{t}$ is the stepsize, and $\epsilon$ prevents division by zero. This step could be further simplified by only taking the instantaneous samples into (19), that is, $\lambda=0$, to yield
@@ IV-D Tuning Hyperparameters
Our algorithm requires tuning several hyperparameters: regularisation constants $\boldsymbol{\mu}_{t}:={[\mu_{1,t},\mu_{2,t},...,\mu_{P,t}]}^{T}$, $\eta_{t}$, $\gamma$ and $\epsilon$; stepsizes $\mathbf{A}_{t}$, $\mathbf{\beta}_{t}$ and $\mathbf{\rho}_{t}$; and the forgetting factor $\lambda$.
@@
IV-E A Note on Convergence
A rigorous convergence analysis for CGP identification is presented in [19]. However, the assumptions for successful convergence are quite restrictive, requiring both specific sparsity structures (Assumption A5 in [19]) and strong stability conditions (Assumptions A4 and A6 in [19]). Attempts to relax these assumptions have had limited success, partly due to inherent biases in the base problem (8). For instance, the $\ell_{1}$-norm regularisation tends to underestimate non-zero elements [36], and the commutator term adds further complexity.
Given these theoretical limitations, we complement our analysis with empirical convergence metrics. Specifically, we evaluate precision, recall, and F1 scores for identifying non-zero elements in $\mathbf{W}$, independent of their exact values.
@@ V-A Data Generation
The AdaCGP algorithm was evaluated across four graph topologies: random (R), Erdős-Rényi (ER) [38], K-Regular (KR), and Stochastic Block Model (SBM) [39], each with $N=50$ nodes. The random graph weights were drawn from $\mathcal{N}(0,1)$, thresholded between $0.3w_{\max}$ and $0.7w_{\max}$ ($w_{\max}$ being the maximum absolute weight), and normalised by $1.5$ times the largest eigenvalue. For the ER graph, edges from $\mathcal{N}(0,1)$ were thresholded between 1.6 and 1.8 in absolute value ($p_{ER}\approx 0.04$), soft thresholded by 1.5 to range between 0.1 and 0.3, and normalised by $1.5$ times the largest eigenvalue. The KR graph assigned 3 nearest-neighbor connections per node with weights from $\mathcal{U}(0.5,1.0)$, normalised by $1.1$ times the largest eigenvalue. The SBM graph comprised $k=10$ equal-sized clusters, with intra-cluster connection probability 0.05 and inter-cluster probabilities uniform in $[0,0.04]$. Edge weights were sampled from a Laplacian distribution with rate $\lambda_{e}=2$, normalised by $1.1$ times the largest eigenvalue.
Following [19], we set model order to $P=3$ and generated sparse filter coefficients $h_{ij}$ ($2\leq i\leq P$, $0\leq j\leq i$) from $2^{i+j}h_{ij}\sim 0.5\mathcal{U}(-1,-0.45)+0.5\mathcal{U}(0.45,1)$, normalised by 1.5 for stability. Graph signals, $\mathbf{x}_{t}\in\mathbb{R}^{N\times 1}$, were then generated recursively from (5) with $\mathbf{x}_{t\leq P}=\mathbf{0}$ and $\mathbf{w}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{N})$. From 10,500 generated samples, we discarded the first 500 to account for transitory effects and retained 10,000 for analysis. After hyperparameter optimisation, this process was repeated 20 times to construct Monte Carlo estimates of our metrics and parameters.
@@ V-B Estimators: AdaCGP Variants and Baseline Models
We evaluated four AdaCGP variants defined by two design choices: which implementation of Algorithm 1 (Path 1, P1, or Path 2, P2), and when to apply Algorithm 2 (after Algorithm 1 reaches steady-state or in alternation at each time step). These variants were: 1) AdaCGP (P1 w/ debias), 2) AdaCGP (P1 w/ alt debias), 3) AdaCGP (P2 w/ debias), and 4) AdaCGP (P2 w/ alt debias).
@@
Steady-state detection was used to perform early stopping of Algorithms 1 and 2. It utilises exponential moving averages of the observable metrics $\text{NMSE}(\mathbf{x}_{\Psi})$ and $\text{NMSE}(\mathbf{x}_{h})$, denoted as $\sigma_{t}^{(1)}$ and $\sigma_{t}^{(2)}$ respectively, where $\sigma_{t}=\alpha\sigma_{t-1}+(1-\alpha)\text{NMSE}_{t}$ with $\alpha=0.995$. Steady-state is reached when the moving average fails to improve by $1\%$ within 250 epochs. For Algorithm 1, we monitored $\sigma_{t}^{(1)}$ convergence to control the switch to Algorithm 2, as at this stage $\hat{\mathbf{h}}$ may not be meaningful due to bias in $\hat{\mathbf{W}}$. For Algorithm 2, $\sigma_{t}^{(2)}$ convergence determines early stopping and defines the terminal epoch. For alternating debiasing, only $\sigma_{t}^{(2)}$ is monitored.
@@
Hyperparameter optimisation for each variant used a random grid search over 10,000 trials, sampling elements of $\boldsymbol{\mu}$ uniformly from (0.001, 1], $\eta$ from (0.005, 0.1] with step 0.005, $\gamma$ from (0.05, 2.0] with step 0.05, and $\lambda$ from (0.80-0.99] with step 0.01. The best parameters were selected by minimising the average of $\text{NMSE}(\mathbf{x}_{h})$ during the final patience epochs of steady-state.
@@
We benchmarked against TISO and TIRSO [24], sparse adaptive VAR models that track time-varying causality graphs online using LMS and RLS approaches, respectively. Unlike our method, these baselines do not model structural dependencies between parameters, like shift-invariance. In these cases, graph topology estimates $\mathbf{W}_{t}$ are derived from the notion of VAR causality [24], where element $(i,j)$ is causal if $\sum_{p=1}^{P}\mathbbm{1}(\boldsymbol{\Psi}_{ij}^{(p)}\neq 0)>0$, with weights assigned to causal edges as $\mathbf{W}_{ij}=\|\boldsymbol{\Psi}_{ij}\|_{2}\cdot\text{causal}_{ij}$. For TISO, the sparsity hyperparameters span (0.025, 1.0] with step 0.025, while TIRSO additionally samples forgetting factors from (0.80-0.99] with step 0.01, optimising for minimal steady-state prediction error $\text{NMSE}(\mathbf{x}_{\Psi})$.
@@
To ensure optimal convergence, all models employed an adaptive step size $\alpha_{t}$ (with $\mathbf{A}_{t}=\alpha_{t}\mathbf{I}_{P}$), computed as
@@
where $\epsilon$ ensures numerical stability. Other stepsizes $\mathbf{\beta}_{t}$ and $\mathbf{\rho}_{t}$ were calculated using the Armijo rule for automatic selection.
TABLE I: Performance comparison of AdaCGP variants and baseline models (TIRSO, TISO) in steady-state for the K-Regular topology. Values show median $\pm$ interquartile range across trials. Best results are shown in bold, and second-best are underlined.
Method NMSE$(\mathbf{W})$ $P_{M}$ $P_{FA}$
TIRSO 0.79 $\pm$ 0.02 0.00 $\pm$ 0.00 1.00 $\pm$ 0.00
TISO 0.78 $\pm$ 0.02 0.00 $\pm$ 0.00 1.00 $\pm$ 0.00
P2 w/ alt. debias 0.51 $\pm$ 0.02 0.02 $\pm$ 0.01 0.61 $\pm$ 0.03
P2 w/ debias 0.33 $\pm$ 0.02 0.01 $\pm$ 0.00 0.53 $\pm$ 0.10
P1 w/ alt. debias 0.21 $\pm$ 0.04 0.01 $\pm$ 0.01 0.00 $\pm$ 0.00
P1 w/ debias 0.14 $\pm$ 0.02 0.00 $\pm$ 0.00 0.01 $\pm$ 0.02
@@ V-C Convergence Across Graph Topologies
Fig. 1 shows the convergence of our metrics for the KR topology as a representative example. All models converge to low non-zero values for NMSE$(\mathbf{x}_{\Psi})$ and NMSE$(\mathbf{x}_{h})$, with Path 1 achieving lower errors than Path 2. The NMSE$(\mathbf{x}_{h})$ for Path 1 showed significant improvement post-debiasing, and achieved near-zero NMSE$(\mathbf{W})$. While alternating debiasing converged to slightly higher values than steady-state debiasing, it demonstrated faster convergence in NMSE$(\mathbf{W})$ for Path 1.
@@
As shown in Table I, the AdaCGP variants consistently outperformed the baselines. The AdaCGP (P1 w/ debias) achieves an 82% reduction in $\text{NMSE}(\mathbf{W})$ compared to TISO ($0.14\pm 0.02$ versus $0.78\pm 0.02$) and a 58% reduction compared to P2 w/ debias ($0.33\pm 0.02$). Notably, Path 1 variants maintained near-zero false alarm rates ($P_{FA}\leq 0.01$) while achieving minimal missing edge rates ($P_{M}\leq 0.01$), in contrast to baseline models’ $P_{FA}$ of 1.00 and Path 2’s $P_{FA}$ exceeding 0.50. The baseline models’ $P_{FA}$ of 1.00 reflected their sparsity mechanisms’ inability to explicitly set elements to zero, while our variable splitting approach overcame this limitation, effectively identifying and removing irrelevant edges for an improved $P_{FA}$ and $P_{M}$ trade-off.
Figure 2: GSO matrices across four graph topologies: (a) Random, (b) Erdős-Rényi, (c) K-Regular, and (d) Stochastic Block Model. For each topology we show: true GSO $\mathbf{W}$ (left), mean estimate $\text{Mean}(\mathbf{\hat{W}})$ over 20 realisations (middle), and most commonly estimated non-zero elements $\text{Mode}(\mathbf{\hat{W}}\neq 0)$ with highlighted true positives and false negatives (right). Results shown for the best-performing AdaCGP variant based on steady-state $\text{NMSE}(\mathbf{x}_{h})$.

### v2

@@ IV-B Updating $\mathbf{W}_{t}$
The next step of our algorithm involves estimating the GSO, $\mathbf{W}_{t}$, from $\hat{\boldsymbol{\Psi}}_{t}$ through either of two paths. Path 2 provides a simple but biased estimate by setting $\hat{\mathbf{W}}_{t}=\hat{\boldsymbol{\Psi}}_{1,t}$. Path 1, which we detail below, optimises (9) to obtain direct estimates of $\mathbf{W}_{t}$ at the expense of computational complexity.
@@
where $\beta_{t}$ is the step-size, and $\mu_{1,t}$ is the sparsity parameter from the first sub-problem. Then, $\hat{\mathbf{W}}_{t}={\hat{\mathbf{W}}}_{+_{t}}-{\hat{\mathbf{W}}}_{-_{t}}$.
@@ IV-C Estimating $\mathbf{h}_{t}$
To fully identify the CGP model in (5), the graph filter coefficients $\hat{\mathbf{h}}_{t}$ remain to be estimated. This phase is optional since the primary objective of this work is to estimate the GSO structure. Notice that, $\hat{\mathbf{W}}_{t}$ must be de-biased before estimating $\hat{\mathbf{h}}_{t}$, due to the heavy regularisation involved. De-biasing is performed by fixing and optimising only the non-zero elements of $\hat{\mathbf{W}}_{t}$ using least squares. Note that reducing the sample size in this way can risk distorting the noise distribution from normality, potentially compromising the minimum MSE criterion for noisy or small datasets [47].
After de-biasing $\hat{\mathbf{W}}_{t}$, the filter coefficients $\hat{\mathbf{h}}_{t}$ are estimated. Unlike zeros in the GSO, which indicate absence of edges and thus have causal significance, sparsity in $\hat{\mathbf{h}}$ serves mainly to simplify the model structure. Given the more relaxed sparsity requirements, GAR-LMS [37] is employed to arrive at the recursive update equation of $\hat{\mathbf{h}}$, given by
@@
where $\mathbf{Y}_{t}$ contains the lagged graph shifted signals as in (11), $\mathbf{C}_{t}\in\mathbb{R}^{M\times M}$ tracks correlations between the shifted signals, $\mathbf{u}_{t}\in\mathbb{R}^{M}$ accumulates cross-correlations with the desired response, $\mathbf{b}_{t}$ is the sparsity-promoting re-weighting vector, $\rho_{t}$ is the step-size, and $\epsilon$ prevents division by zero. This step could be further simplified by only taking the instantaneous samples into (19), that is, $\lambda=0$, to yield
@@ IV-D Tuning Hyperparameters
Our algorithm requires tuning several hyper-parameters: regularisation constants $\boldsymbol{\mu}_{t}:={[\mu_{1,t},\mu_{2,t},...,\mu_{P,t}]}^{T}$, $\eta_{t}$, $\gamma$ and $\epsilon$; step-sizes $\mathbf{A}_{t}$, $\mathbf{\beta}_{t}$ and $\mathbf{\rho}_{t}$; and the forgetting factor $\lambda$.
@@
IV-E Discussion on Convergence
A rigorous convergence analysis for batch CGP identification using the same alternating steps of convex regularised least squares sub-problems was presented in [28]. Their analysis bounded the average excess prediction risk, defined as the difference between prediction error using the estimated and true parameters. This error scales as $O(\sqrt{\log N/K})$ for estimating the GSO, $\mathbf{W}$, and as $O(\sqrt{\log N/K^{2(\nu-\beta)}})$ for estimating the filter coefficients, $\mathbf{h}$, where $N$ is the network size, $K$ is the number of time observations, and $\nu,\beta$ are parameters used in their high-probability bounds that satisfy $0<\beta<\nu<1/2$.
AdaCGP extends this framework through variable splitting as expressed in (12) and (13), while maintaining the same core optimisation structure as [28]. The variable splitting approach reformulates the $\ell_{1}$-regularised problem as a non-negativity constrained optimisation [46], aligning with two-metric projection methods [50]. For such projection methods, [50] established that for non-critical points, there exists a step size that guarantees descent, and that every limit point of the generated sequence is a critical point. When applied to our non-negativity constraints resulting from variable splitting, these results provide theoretical support for our algorithm’s convergence properties, with projected gradient theory suggesting similar convergence behaviour under standard convexity and step size conditions. The effectiveness of this technique for $\ell_{1}$-regularised problems has been empirically demonstrated by [46], showing consistent improvements in convergence speed and computational efficiency.
The relationship between batch and online optimisation algorithms is well-established. For adaptive filtering, LMS converges to the Wiener-Hopf solution under small step sizes and stationarity [32], and RLS variants achieve faster convergence rates [33, 34]. Both approaches solve the same fundamental optimisation problem with different processing mechanisms: batch methods process all data simultaneously with multiple passes, while online methods process samples sequentially with a forgetting factor $\lambda$. This suggests the convergence properties established for the batch CGP algorithm [28] and variable splitting [50] should extend to our adaptive setting, given stationarity, appropriate forgetting factor, and sufficient iterations.
Despite these theoretical guarantees, practical limitations remain. The theoretical framework in [28] relies on key assumptions regarding sparsity structures and stability conditions that may not hold in many real-world networks. Additionally, $\ell_{1}$-regularisation introduces shrinkage bias [48], and the commutativity term in (8) adds further complexity. Given these theoretical limitations, we complement our analysis with empirical convergence metrics, evaluating AdaCGP’s ability to correctly identify the graph topology through precision, recall, and F1 scores for non-zero elements in $\mathbf{W}$. This approach recognises that identifying the presence or absence of connections is often more important than precisely estimating their weights.
@@ V-A Data Generation
The AdaCGP algorithm was evaluated across four graph topologies: Random (R), Erdős-Rényi (ER) [51], K-Regular (KR), and Stochastic Block Model (SBM) [52], each with $N=50$ nodes. The random graph weights were drawn from $\mathcal{N}(0,1)$, thresholded between $0.3w_{\max}$ and $0.7w_{\max}$ ($w_{\max}$ being the maximum absolute weight), and normalised by $1.5$ times the largest eigenvalue. For the ER graph, edges from $\mathcal{N}(0,1)$ were thresholded between 1.6 and 1.8 in absolute value ($p_{ER}\approx 0.04$), soft thresholded by 1.5 to range between 0.1 and 0.3, and normalised by $1.5$ times the largest eigenvalue. The KR graph assigned 3 nearest-neighbour connections per node with weights from $\mathcal{U}(0.5,1.0)$, normalised by $1.1$ times the largest eigenvalue. The SBM graph comprised $k=10$ equal-sized clusters, with intra-cluster connection probability 0.05 and inter-cluster probabilities uniform in $[0,0.04]$. Edge weights were sampled from a Laplacian distribution with rate $\lambda_{e}=2$, normalised by $1.1$ times the largest eigenvalue. Following [28], different normalisation factors (1.1 for KR/SBM, 1.5 for Random/ER) ensure stability while preserving graph-specific properties: structured graphs require less conservative scaling than those with variable spectral properties. These factors work with filter coefficient normalisation to maintain stability in signal generation from (5).
Following [28], we set model order to $P=3$ and generated sparse filter coefficients $h_{ij}$ ($2\leq i\leq P$, $0\leq j\leq i$) from $2^{i+j}h_{ij}\sim 0.5\mathcal{U}(-1,-0.45)+0.5\mathcal{U}(0.45,1)$, normalised by 1.5 for stability. Graph signals, $\mathbf{x}_{t}\in\mathbb{R}^{N\times 1}$, were then generated recursively from (5) with $\mathbf{x}_{t\leq P}=\mathbf{0}$ and $\mathbf{w}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{N})$. From 11,000 generated samples, we discarded the first 1000 to account for transitory effects and retained 10,000 for analysis. After hyper-parameter optimisation, this process was repeated 5 times to construct Monte Carlo estimates of our metrics and parameters.
@@ V-B Estimators: AdaCGP Variants and Baseline Models
We evaluated four AdaCGP variants defined by two design choices: which implementation of Algorithm 1 to use (Path 1 or Path 2), and when to apply Algorithm 2 (after Algorithm 1 reaches steady-state versus in alternation at each time step). These variants were: 1) AdaCGP (P1 + debias), 2) AdaCGP (P1 + alt debias), 3) AdaCGP (P2 + debias), and 4) AdaCGP (P2 + alt debias). For all variants, we used simplified instantaneous updates ($\lambda=0$ in (19)) for $\hat{\mathbf{h}}$ to prioritise computational efficiency, as our main goal was accurate GSO estimation rather than optimal $\text{NMSE}(\mathbf{x}_{h})$ values.
@@
Steady-state detection was used to perform early stopping of Algorithms 1 and 2. It utilises exponential moving averages of the observable metrics $\text{NMSE}(\mathbf{x}_{\Psi})$ and $\text{NMSE}(\mathbf{x}_{h})$, denoted as $\sigma_{t}^{(1)}$ and $\sigma_{t}^{(2)}$ respectively, where $\sigma_{t}=\alpha\sigma_{t-1}+(1-\alpha)\text{NMSE}_{t}$ with $\alpha=0.995$. Steady-state is reached when the moving average fails to improve by $1\%$ within 500 epochs. For Algorithm 1, we monitored $\sigma_{t}^{(1)}$ convergence to control the switch to Algorithm 2, as at this stage $\hat{\mathbf{h}}$ may not be meaningful due to bias in $\hat{\mathbf{W}}$. For Algorithm 2, $\sigma_{t}^{(2)}$ convergence determines early stopping and defines the terminal epoch. For alternating debiasing, only $\sigma_{t}^{(2)}$ is monitored.
@@
Hyper-parameter optimisation for each variant used a random grid search over 20,000 trials, sampling elements of $\boldsymbol{\mu}$ uniformly from (0.001, 1], $\eta$ from (0.005, 0.1] with step 0.005, $\gamma$ from (0.05, 2.0] with step 0.05, and $\lambda$ from (0.80-0.99] with step 0.01. The best parameters were selected by minimising the average of $\text{NMSE}(\mathbf{x}_{h})$ during the final patience epochs of steady-state.
@@
We benchmarked against several established baselines, both adaptive and batch methods. For adaptive methods, we included TISO and TIRSO [35], sparse adaptive VAR models that track time-varying causality graphs online using LMS and RLS approaches, respectively. Unlike our method, these baselines do not model structural dependencies between parameters, like shift-invariance. In these cases, graph topology estimates $\hat{\mathbf{W}}_{t}$ are derived from the notion of VAR causality [35], where element $(i,j)$ is causal if $\sum_{p=1}^{P}\mathbbm{1}(\boldsymbol{\Psi}_{ij}^{(p)}\neq 0)>0$, with weights assigned to causal edges as $\hat{\mathbf{W}}_{ij}=\|\hat{\boldsymbol{\Psi}}_{ij}\|_{2}\cdot\text{causal}_{ij}$. For TISO, the sparsity hyper-parameters span (0.025, 1.0] with step 0.025, while TIRSO additionally samples forgetting factors from (0.80-0.99] with step 0.01, optimising for minimal steady-state prediction error $\text{NMSE}(\mathbf{x}_{\Psi})$.
We also included SD-SEM [36], an adaptive method that tracks evolving topologies using SEMs, which model observations as linear combinations of endogenous variables with sparsity assumptions. The hyper-parameter optimisation considered forgetting factors (0.80-0.99] and sparsity penalties with values 1, 5, 10, 15, 20, and 25.
For batch methods, we included the graphical LASSO (GLasso) [15], which estimates a sparse precision matrix from multivariate normal observations using the Markov property, with sparsity parameter ranging from (0.10-1.0] with step 0.10. We also evaluated GL-SigRep [23], which learns a graph Laplacian matrix under signal smoothness assumptions. We applied GL-SigRep per time-instant, and the hyper-parameter optimisation considered smoothness regularisation values $10^{-3}$, $10^{-2.5}$, $10^{-2}$, $10^{-1.5}$, $10^{-1}$ and sparsity strengths $10^{-2}$, $10^{-1.5}$, $10^{-1}$, $^{-0.5}, 1001$$^{0}$. Finally, we included VAR and VAR+Granger [20] models with order $P=3$, where Granger causality determines directed edges through statistical testing of VAR model coefficients.
All batch methods, except GL-SigRep, were trained with all data available up to time $t$. During training, the respective objective function values were monitored for convergence, and metrics were computed on the final patience epochs of steady-state. Except for VAR+Granger, where causality tests were performed on the last 50 iterations for computational efficiency.
@@
To ensure optimal convergence, all adaptive autoregressive models employed an adaptive step size $\alpha_{t}$ (with $\mathbf{A}_{t}=\alpha_{t}\mathbf{I}_{P}$), computed as
@@
where $\epsilon$ ensures numerical stability. For SD-SEM, which is not autoregressive, the default step size of $1/\lambda_{\text{max}}$ was used. Other step-sizes $\mathbf{\beta}_{t}$ and $\mathbf{\rho}_{t}$ were calculated using the Armijo rule for automatic selection.
TABLE I: Comparison of graph topology estimation algorithms for $N=50$. Lower values are better for all metrics.
Best results are in bold, second best are underlined.
Random Erdos-Renyi
Algorithm $_{\text{FC}} NMSE$$(\mathbf{W})$ PM PFA $_{\text{FC}} NMSE$$(\mathbf{W})$ PM PFA
Batch
[…]
Adaptive
[…]
K-Regular Stochastic Block Model
Algorithm $_{\text{FC}} NMSE$$(\mathbf{W})$ PM PFA $_{\text{FC}} NMSE$$(\mathbf{W})$ PM PFA
Batch
[…]
Adaptive
[…]
@@ V-C Convergence Across Graph Topologies
[…]
@@
As shown in Table I, the AdaCGP variants consistently outperformed the baselines across all graph topologies. To facilitate comparison across algorithms, we include $\text{NMSE}_{\text{FC}}$ in Table I, representing the forecast error of each model (if available), which equals $\text{NMSE}(\mathbf{x}_{\mathbf{h}})$ for AdaCGP variants.
Specifically, AdaCGP (P1 + debias) achieves the lowest or second-lowest $\text{NMSE}(\mathbf{W})$ across all topologies, with improvements in excess of $94\%$, $98\%$, $83\%$, and $93\%$ for Random, Erdos-Renyi, K-Regular, and SBM topologies, respectively, compared to baseline models. Among adaptive algorithms, TISO and TIRSO fail to enforce sparsity ($P_{FA}=1.00$ across all topologies), while SD-SEM sets some edges to zero but with consistently high error rates. In contrast, AdaCGP Path 1 variants maintain near-zero false alarm rates ($P_{FA}\leq 0.08$) while achieving minimal missing edge rates ($P_{M}\leq 0.12$) across all topologies. These results highlight that baseline methods struggle with their sparsity mechanisms’ inability to explicitly set elements to zero, while our variable splitting approach overcomes this limitation, effectively identifying and removing irrelevant edges for an improved $P_{FA}$ and $P_{M}$ trade-off. While for batch methods, GLasso achieves lower $P_{FA}$ values but at the expense of significantly higher $P_{M}$.
The SBM topology presents a particularly challenging case due to its high sparsity ($\approx 1\%$ non-zero elements). Here, AdaCGP (P1 + debias) is surpassed only by VAR+Granger in $P_{M}$ ($0.12\pm 0.08$ vs. $0.08\pm 0.04$) and $P_{FA}$ ($0.08\pm 0.29$ vs. $0.05\pm 0.00$). The higher variance in $P_{FA}$ results from debiasing being applied upon convergence when facing extremely sparse topologies, where minor edge detection errors significantly impact metrics. However, despite methodological differences between VAR+Granger (sparsity obtained by offline statistical testing) and AdaCGP (sparsity obtained online), AdaCGP (P1 + debias) achieves a $\approx$$39\times$ reduction in $\text{NMSE}(\mathbf{W})$ ($0.09\pm 0.17$ vs. $3.50\pm 0.03$).
It is important to note that $\text{NMSE}(\mathbf{x}_{h})$ and $\text{NMSE}(\mathbf{W})$ evaluate different aspects of model performance (prediction accuracy vs. topology estimation accuracy). Due to the structure of our alternating optimisation framework in (8)-(10), minimising $\text{NMSE}(\mathbf{x}_{h})$ often leads to improved $\text{NMSE}(\mathbf{W})$, while the converse is not necessarily true.
[…]
V-D Identification of the GSO Topology
[…]